OpenStudy (anonymous):

the function given by f(x)=2(x+10) can be used to convert dress sizes x in the US to dress sizes f in Italy. For what dress sizes in the US will dress sizes in Italy be between 32 and 46?

OpenStudy (anonymous):

no worries your still the man

jimthompson5910 (jim_thompson5910):

so f(x) = 32 and f(x) = 46

jimthompson5910 (jim_thompson5910):

32 <= f(x) <= 46 32 <= 2(x+10) <= 46 32 <= 2x+20 <= 46 32-20 <= 2x <= 46-20 12 <= 2x <= 26 12/2 <= x <= 26/2 6 <= x <= 13 So a size 6 in the US corresponds to a size 32 in Italy Also a size 13 in the US corresponds to a size 46 in Italy